### v2

# Nonlocal characteristics of two-qubit gates and their argand diagrams

Abstract
In this paper, we show the usefulness of the chords present in the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates to study their nonlocal characteristics. We discuss the criteria for perfect entanglers to transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states. Perfect entanglers with a chord passing through origin can do such a transformation. In the Weyl chamber, we identify the regions of perfect entanglers with at least one chord passing through origin. We also provide the conditions for a perfect entangler without any chord passing through origin to transform a pair of orthonormal product states into orthonormal maximally entangled states. Finally, we show that similar to entangling power, gate typicality can also be described using the chords present in the argand diagram. For each chord describing the entangling power, there exists a chord describing the gate typicality. We show the geometrical relation between the two sets of chords.
preprint: APS/123-QED
I Introduction
Understanding the geometry and nonlocal characteristics of two-qubit gates is essential as their role in quantum computation is vital [1, 2, 3]. Over two decades, two-qubit gates have been studied vastly. Nonlocal characteristics of two-qubit gates are invariant under local operations. A complete set of local invariants of two-qubit gates was obtained in [4]. The geometry of local equivalence classes of two-qubit gates was studied in detail in [5]. A measure for operator entanglement was introduced in [6]. Nonlocal characteristics of two-qubit gates were considered as resources for doing quantum information processing and their quantification were studied in [7]. Entangling power, a quantity to measure the ability of two-qubit gates to generate entanglement, was defined in [8]; its expression in terms of Cartan co-ordinates [9] and local invariants [10] were derived. Gate typicality as complementary to entangling power was introduced [11] and its properties for two-qubit gates were studied [12].
Despite many studies, still there remain unexplored ways to understand the nonlocal characteristics of two-qubit gates. Recently, the chords present in the argand diagram of squared eigenvalues of nonlocal parts of two-qubit gates were shown to describe the ability of two-qubit gates to generate entangled states [13]. This argand diagram was studied in regard to the condition for perfect entanglers [5, 4], operational discrimination of the nonlocal part of a two-qubit gate from the nonlocal part of its adjoint [14], and simulation of perfect entanglers [15]. Perfect entanglers are two-qubit gates that can transform some product state into a maximally entangled state. For perfect entanglers, the convex hull of squared eigenvalues of their nonlocal part contains zero [5, 4]. Among perfect entanglers, special perfect entanglers (SPEs) can transform orthonormal product basis into orthonormal maximally entangled basis [9]. Recently, it was shown that perfect entanglers with a chord passing through origin in their argand diagram can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states [13]. Now, there arises a question that can a perfect entangler without any chords passing through origin transforms a pair of orthonormal product states into a pair of orthonormal maximally entangled states?
In this paper, we provide answer to this question. First, we identify the regions of perfect entanglers with at least one chord passing through origin. Next, we derive two sets of orthogonal conditions between a pair of product states obtained from the argand diagram of perfect entanglers without any chord passing through origin. We express them in terms of eigenvalues of generators of perfect entanglers. Satisfying one of these conditions is both necessary and sufficient for a perfect entangler without any chord passing through origin to transform a pair of orthonormal product states into maximally entangled states. Finally, we show that in addition to entangling power, the chords present in the argand diagram can also be used to quantify the gate typicality and linear entropy [6] of two-qubit gates.
II Background
A two-qubit gate $U\in\textnormal{SU}(4)$ can be decomposed as follows [5].
$U=L_{1}U_{d}(c_{x},c_{y},c_{z})L_{2},$ (1)
where $L_{1},~{}L_{2}\in\textnormal{SU}(2)\otimes\textnormal{SU}(2)$ are local parts of $U$ and $U_{d}(c_{x},c_{y},c_{z})$ is the nonlocal part of $U$. Two-qubit gates having the same nonlocal part but differing in the local parts form a local equivalence class and each local equivalence class of two-qubit gates is geometrically represented as a point of tetrahedron shown in Fig. 1.
Nonlocal characteristics of $U$ is determined by its nonlocal part which can be written as
$U_{d}(c_{x},c_{y},c_{z})=e^{iH/2}$ (2)
with
$H=\sum_{j=x,y,z}c_{j}\left(\sigma_{j}\otimes\sigma_{j}\right),$ (3)
where $c_{x},c_{y},c_{z}$ are called Cartan co-ordinates. Eigenvalues of $H$ are given by
$h_{1}=c_{x}-c_{y}+c_{z},$
$h_{2}=c_{x}+c_{y}-c_{z},$
$h_{3}=-c_{x}-c_{y}-c_{z},$
and
$h_{4}=-c_{x}+c_{y}+c_{z}.$ (4)
Figure 1: Geometry of local equivalence classes of two-qubit gates. Region coloured in blue is the region of perfect entanglers. The line in red colour represent special perfect entanglers.
Eigen decomposition of the nonlocal part of $U$ can be written as shown below.
$U_{d}(c_{1},c_{2},c_{3})=\sum_{j=1}^{4}e^{ih_{j}/2}|\Psi_{j}\rangle\langle\Psi_{j}|,$ (5)
where the column vectors $|\Psi_{j}\rangle$ form the eigen basis of $U_{d}$. It is also called as magic basis [5] and their expressions are given below.
$|\Psi_{1}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}1\\ 0\\ 0\\ 1\end{tabular}\right],~{}|\Psi_{2}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}0\\ i\\ i\\ 0\end{tabular}\right],$ (14)
$|\Psi_{3}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}0\\ 1\\ -1\\ 0\end{tabular}\right],~{}|\Psi_{4}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}i\\ 0\\ 0\\ -i\end{tabular}\right].$ (23)
The squared eigenvalues of $U_{d}$ are on a circle of unit radius in the complex plane and they are pairwise connected by six chords forming a quadrilateral as shown in Fig. 2. We refer to this argand diagram of squared eigenvalues of $U_{d}$ as the argand diagram of $U$.
The Cartan co-ordinates of perfect entanglers which can transform a product state into a maximally entangled state satisfy the following conditions [16].
$c_{x}+c_{y}\geq\frac{\pi}{2},$
$c_{y}\pm c_{z}\leq\frac{\pi}{2}.$ (24)
The region of perfect entanglers (coloured in blue) is shown in Fig. 1. Special perfect entanglers which are perfect entanglers with maximum entangling power [9] are represented by the red colour line in Fig. 1. For perfect entanglers, the convex hull of squared eigenvalues of $U_{d}$ encloses zero. The argand diagram of a typical perfect entangler (non perfect entangler) is shown in Fig. 2a (Fig. 2b).
Figure 2: Argand diagram of squared eigenvalues of (a) a perfect entangler, and (b) a non perfect entangler.
III Perfect entanglers with at least a chord passing through origin
In this section, we identify the regions of perfect entanglers with at least one chord passing through origin. The chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$ passes through origin only if $|h_{j}-h_{k}|=\pi$. It can be verified that the equations $|h_{j}-h_{j+1}|=\pi$ with $j=1,2,3$ describe the three boundary planes separating perfect entanglers from non perfect entanglers inside the Weyl chamber [Eq. 24]. These three planes are shown in Fig. 3a. Similarly, the equations $|h_{j}-h_{j+2}|=\pi$ with $j=1,2$ describe the planes $c_{x}\pm c_{z}=\pi/2$ [Fig. 3b]. These two planes together form one of the reflecting plane, $c_{x}+\textnormal{sign}(c_{z})c_{z}=\pi/2$, describing the mirror operation [17]. The equation $|h_{1}-h_{4}|=\pi$ which can be rewritten as $c_{x}-c_{y}=\pi/2$ is satisfied only by the point representing CNOT equivalence class. Thus only the gates represented by the five planes shown in Fig. 3a and Fig. 3b have at least one chord passing through origin and hence they can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states.
All the gates represented by a specific plane have unique argand diagram with a specific chord passing through origin. However, there are gates present in more than one plane. SPEs are at the intersection of $|h_{j}-h_{j+2}|=\pi~{}(j=1,2)$ planes. Hence, for SPEs, two chords connecting diametrically opposite points $(e^{ih_{j}}~{}\textnormal{and}~{}e^{ih_{j+2}}~{}\textnormal{with}~{}j=1,2)$ pass through origin and they can transform orthonormal product basis into orthonormal maximally entangled basis [13, 9]. Among SPEs, the point representing iSWAP equivalence class is also present in $|h_{j}-h_{j+1}|=\pi~{}\textnormal{with}~{}j=1,3$ planes. Similarly, the point representing CNOT equivalence class also satify the conditions, $|h_{2}-h_{3}|=\pi$ and $|h_{1}-h_{4}|=\pi$. Hence, these two equivalence classes have argand diagrams with four chords passing through origin [13].
Similar to SPEs, there are local equivalence classes at the intersection of two planes: between the pairs of planes $\{|h_{3}-h_{4}|=\pi,|h_{1}-h_{3}|=\pi\}$, $\{|h_{2}-h_{3}|=\pi,|h_{1}-h_{3}|=\pi\}$, $\{|h_{1}-h_{2}|=\pi,|h_{2}-h_{4}|=\pi\}$, and $\{|h_{2}-h_{3}|=\pi,|h_{2}-h_{4}|=\pi\}$. However, for these equivalence classes, the two chords connecting diametrically opposite points do not pass through origin but, only one of them passes through origin. Hence these equivalence classes cannot transform orthonormal product basis into orthonormal maximally entangled basis.
The point representing $\sqrt{\textnormal{SWAP}}$ equivalence class is common to $|h_{3}-h_{4}|=\pi$, $|h_{1}-h_{3}|=\pi$, and $|h_{2}-h_{3}|=\pi$ planes. Hence, in the argand diagram of $\sqrt{\textnormal{SWAP}}$ equivalence class, three chords from the point $e^{ih_{3}}$ pass through origin. It implies that the remaining three points $e^{ih_{1}}$, $e^{ih_{2}}$, and $e^{ih_{4}}$ coincide with each other and the argand diagram has only three chords. Similarly, it can be verified that the argand diagram of $\sqrt{\textnormal{SWAP}}^{\dagger}$ equivalence class has only three chords from the point $e^{ih_{2}}$ passing through origin. Thus for the gates belonging to these two equivalence classes there exist three different pairs of orthonormal product states that can be transformed into pairs of orthonormal maximally entangled states.
Figure 3: Regions of perfect entanglers with at least one chord passing through origin. In the subfigure (a) the planes coloured in magenta, green, and yellow are $|h_{1}-h_{2}|=\pi$, $|h_{2}-h_{3}|=\pi$, and $|h_{3}-h_{4}|=\pi$ planes respectively. The violet, purple, and orange lines (excluding the end points) represent perfect entanglers without any chord passing through origin but capable of transforming a pair of orthonormal product states into a maximally entangled states. These perfect entanglers satisfy the conditions obtained in case 2. In the subfigure (b), the planes coloured in red and blue are $|h_{1}-h_{3}|=\pi$ and $|h_{2}-h_{4}|=\pi$ planes respectively.
IV Perfect entanglers without any chord passing through origin
In this section, we consider the perfect entanglers without any chord passing through origin. The argand diagram of such a perfect entangler is shown in Fig. 2a. The origin can be expressed as convex combination of at most three points [18]. In Fig. 2a, origin is contained in two triangles: the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$, and the triangle with vertices $e^{ih_{k}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. Hence, there exist two sets of weights or co-ordinates such that
$|\phi_{j}|^{2}e^{ih_{j}}+|\phi_{l}|^{2}e^{ih_{l}}+|\phi_{m}|^{2}e^{ih_{m}}=0,$ (25)
and
$|\phi^{\prime}_{k}|^{2}e^{ih_{k}}+|\phi^{\prime}_{l}|^{2}e^{ih_{l}}+|\phi^{\prime}_{m}|^{2}e^{ih_{m}}=0,$ (26)
with $|\phi_{j}|^{2}+|\phi_{l}|^{2}+|\phi_{m}|^{2}=1$ and $|\phi^{\prime}_{k}|^{2}+|\phi^{\prime}_{l}|^{2}+|\phi^{\prime}_{m}|^{2}=1$. From Eqs. 25 and 26, following two product states can be constructed [5]
$|\Phi_{1}\rangle=|\phi_{j}|e^{-i(a\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi_{m}|e^{-i(c\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (27)
and
$|\Phi_{2}\rangle=|\phi^{\prime}_{k}|e^{-i(d\pi+h_{k}/2}|\Psi_{k}\rangle+|\phi^{\prime}_{l}|e^{-i(f\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi^{\prime}_{m}|e^{-i(g\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (28)
for some integers $a,b,c,d,f,$ and $g$. Now we consider the following cases.
Case 1: First, we consider the condition for the two product states belonging to two different triangles (Eqs. 27 and 28) to be orthogonal. This condition can be written as follows.
$|\phi_{l}||\phi^{\prime}_{l}|e^{i(f-b)\pi}=-|\phi_{m}||\phi^{\prime}_{m}|e^{i(g-c)\pi}.$ (29)
Choosing $(f-b)$ as even (odd) and $(g-c)$ as odd (even), above equation can be rearranged as follows.
$\frac{|\phi_{l}|}{|\phi_{m}|}=\frac{|\phi^{\prime}_{m}|}{|\phi^{\prime}_{l}|}.$ (30)
This is satisfied when the following condition is satisfied.
$\frac{|\phi_{l}|^{2}}{|\phi_{m}|^{2}}=\frac{|\phi^{\prime}_{m}|^{2}}{|\phi^{\prime}_{l}|^{2}}.$ (31)
The weight $|\phi_{l}|^{2}$ can be written as the ratio of the area of subtriangle with vertices $[\cos(h_{j}),\sin(h_{j})]$, $[0,0]$, and $[\cos(h_{m}),\sin(h_{m})]$ to the area of triangle with vertices $[\cos(h_{j}),\sin(h_{j})]$, $[\cos(h_{l}),\sin(h_{l})]$, and $[\cos(h_{m}),\sin(h_{m})]$ [19]. Similarly, other weights can also be expressed and the condition for the product states given in Eqs. 27 and 28 to be orthogonal can be expressed in terms of the eigenvalues of $H$ as follows.
$\frac{|\sin(h_{j}-h_{m})|}{|\sin(h_{l}-h_{j})|}=\frac{|\sin(h_{l}-h_{k})|}{|\sin(h_{k}-h_{m})|}.$ (32)
Thus, if the eigenvalues of $H$ satisfy the above condition, then it is possible to construct a pair of orthonormal product states that can be converted into a pair of orthonormal maximally entangled states by the perfect entangler $e^{iH/2}$ with the argand diagram shown in Fig. 2a.
Case 2: Now, we consider the case where $e^{ih_{j}}$ coincides with $e^{ih_{k}}$ in Fig. 2a. In this case, we have $|\phi^{\prime}_{k}|=|\phi_{j}|$, $|\phi^{\prime}_{l}|=|\phi_{l}|$, and $|\phi^{\prime}_{m}|=|\phi_{m}|$ in Eqs. 26 and 28. Choosing $(f-b)$ as even (odd) and $(g-c)$ as odd (even), the condition for the two product states (Eqs. 27 and 28) to be orthogonal becomes
$|\phi_{l}|^{2}=|\phi_{m}|^{2}.$ (33)
This condition can be expressed in terms of eigenvalues of $H$ as follows.
$\frac{|\sin(h_{j}-h_{m})|}{|\sin(h_{l}-h_{j})|}=1.$ (34)
Case 3: Now, we derive the condition for the existence of a pair of orthonormal product states with respect to a triangle enclosing the origin. For example, in Fig. 2a, we consider the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. A product state corresponding to this triangle is given in Eq. 27. We construct another product state with respect to the same triangle, for some integers $a^{\prime},b^{\prime},~{}\textnormal{and}~{}c^{\prime}$, as shown below.
$|\Phi^{\prime}_{1}\rangle=|\phi_{j}|e^{-i(a^{\prime}\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b^{\prime}\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi_{m}|e^{-i(c^{\prime}\pi+h_{m}/2)}|\Psi_{m}\rangle.$ (35)
These two product states, $|\Phi_{1}\rangle$ and $|\Phi^{\prime}_{1}\rangle$, can be converted into maximally entangled states by $U_{d}(c_{1},c_{2},c_{3})$. These two states to be orthogonal the following condition needs to be satisfied.
$|\phi_{j}|^{2}e^{i(a^{\prime}-a)\pi}+|\phi_{l}|^{2}e^{i(b^{\prime}-b)\pi}+|\phi_{m}|^{2}e^{i(c^{\prime}-c)\pi}=0.$ (36)
This condition will be satisfied only if either one of $(a^{\prime}-a)$, $(b^{\prime}-b)$, and $(c^{\prime}-c)$ is even (odd) and other two are odd (even). If $(a^{\prime}-a)$ is chosen as even (odd) and the other two as odd (even), then we get
$|\phi_{j}|^{2}=|\phi_{l}|^{2}+|\phi_{m}|^{2}.$ (37)
Along with the condition $|\phi_{j}|^{2}+|\phi_{l}|^{2}+|\phi_{m}|^{2}=1$, it can be found that $|\phi_{j}|^{2}=1/2$. This condition can be written in terms of eigenvalues of $H$ as follows.
$\frac{2|\sin(h_{m}-h_{l})|}{|\sin(h_{l}-h_{j})+\sin(h_{j}-h_{m})+\sin(h_{m}-h_{l})|}=1.$ (38)
This condition will be satisfied only if the following equation is satisfied.
$\sin(h_{l}-h_{j})+\sin(h_{j}-h_{m})=\sin(h_{m}-h_{l})$ (39)
Taking $h_{l}-h_{j}=\theta_{1}$ and $h_{j}-h_{m}=\theta_{2}$ for some $\theta_{1}$ and $\theta_{2}$ which could be either positive or negative, we get
$\sin(\theta_{1})+\sin(\theta_{2})=-\sin(\theta_{1}+\theta_{2})$ (40)
This condition is satisfied when either $\theta_{1}=\pm\pi$ or $\theta_{2}=\pm\pi$. This is also satisfied when $\theta_{1}=\pm\pi$ and $\theta_{2}=\pm\pi$, but in this case, $e^{ih_{l}}$ and $e^{ih_{m}}$ coincide with each other. Thus, the weight $|\phi_{j}|^{2}$ to be 0.5, one of the chords from $e^{ih_{j}}$ should pass through origin. Thus, to construct a pair of orthogonal product states (which can be converted into maximally entangled states) with respect to a triangle containing origin, the origin should be at one of the edges of the triangle. So, this case covers only the perfect entanglers with a chord passing through origin.
Now, we consider two product states, $|\xi\rangle=\sum_{j}\epsilon_{j}|\Psi_{j}\rangle$ and $|\xi^{\prime}\rangle=\sum_{j}\epsilon^{\prime}_{j}|\Psi_{j}\rangle$ $(j=1,2,3,4)$ in the magic basis with the conditions $\sum_{j}|\epsilon_{j}|^{2}=\sum_{j}|\epsilon^{\prime}_{j}|^{2}=1$ and $\sum_{j}\epsilon_{j}^{2}=\sum_{j}\epsilon^{\prime 2}_{j}=0$ [5]. If these two product states are transformed into maximally entangled state by $U_{d}$ [Eq. 5] with $h_{j}-h_{k}\neq\pi$ for all $j\neq k$, then we should have (up to a global phase) $\epsilon_{j}=\pm|\epsilon_{j}|e^{-ih_{j}/2}$ and $\epsilon^{\prime}_{j}=\pm|\epsilon^{\prime}_{j}|e^{-ih_{j}/2}$ for all $j$ [5]. In addition, we should have $\sum_{j}|\epsilon_{j}|^{2}e^{-ih_{j}}=\sum_{j}|\epsilon^{\prime}_{j}|^{2}e^{-ih_{j}}=0$ [5]. Since, the squared eigenvalues of $U_{d}$ ($e^{-ih_{j}}$) exist on a plane, one of the weights should be zero in both summations [18]. It can be verified that if the product states, $|\xi\rangle$ and $|\xi^{\prime}\rangle$, are orthogonal then the argand diagram of $U_{d}$ belongs to the kinds which satisfy one of the conditions obtained in cases 1 and 2.
Case 2 describes the perfect entanglers on $c_{1}=c_{2}$ and $c_{2}=\pm c_{3}$ planes of Weyl chamber. On $c_{1}=c_{2}$ plane, the perfect entanglers are contained in the triangular region with vertices $[\pi/4,\pi/4,\pm\pi/4]$ and $[\pi/2,\pi/2,0]$ [Fig. 3a]. All three edges of the triangular region represent perfect entanglers with at least a chord passing through origin. The interior is a region of perfect entanglers with $h_{1}=h_{4}$ and hence only the condition $|\phi_{2}|^{2}=|\phi_{3}|^{2}$ is applicable. It can be verified that only $U_{d}(c_{1},c_{1},0)$ with $c_{1}\in(\pi/4,\pi/2)$ satisfy the condition $|\phi_{2}|^{2}=|\phi_{3}|^{2}$. Thus, in this region of perfect entanglers without any chord passing through origin, only the perfect entanglers represented by the open line segment between the points $[\pi/4,\pi/4,0]$ and $[\pi/2,\pi/2,0]$ (violet coloured line in Fig. 3a) can transform a pair of orthonormal product states into maximally entangled states and the perfect entanglers represented by other interior points cannot do such a transformation. Similarly, it can be shown that on $c_{2}=\pm c_{3}$ planes only $U_{d}(\pi/2,c_{2},\pm c_{2})$ with $c_{2}\in(0,\pi/4)$ (purple and orange coloured lines in Fig. 3a) can transform a pair of orthonormal product states into maximally entangled states. Conditions obtained in case 1 are applicable to the perfect entanglers represented by the points other than those existing on the five planes discussed in the previous section and $c_{1}=c_{2}$, $c_{2}=\pm c_{3}$ planes of perfect entanglers polyhedron [Fig. 1].
V Quantification of nonlocal characteristics using chords
Like two-qubit states, two-qubit gates also possess entanglement and can be quantified using linear entropy [6, 20]. Two-qubit gates represented by $c_{z}$ axis of the Weyl chamber have maximum value of linear entropy [21]. Entangling power is another nonlocal measure of two-qubit gates that quantifies the average entanglement generated over uniform distribution of product states [8, 9]. Gate typicality is a nonlocal measure which is complementary to entangling power [11, 12]. Both entangling power $(e_{p})$ and gate typicality $(g_{t})$ together describe the linear entropy $(L)$ of two-qubit gates. The relationship between them can be written as follows.
$L=\frac{3}{8}\left[3e_{p}+g_{t}\right].$ (41)
Recently, the entangling power of two-qubit gates was shown to be proportional to the mean squared length of the chords in the argand diagram of two-qubit gates [13]. Entangling power can be written in terms of squared chord lengths as follows.
$e_{p}=\frac{1}{72}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{ih_{k}}|^{2}.$ (42)
Figure 4: Diagram explaining the geometrical relation between $|e^{ih_{j}}-e^{ih_{k}}|$ and $|e^{ih_{j}}-e^{-ih_{k}}|$.
Entangling power of a two-qubit gate $U$ can also be expressed as symmetric combination of $L(U)$ and $L(US)-L(S)$ as follows [20, 21, 11].
$e_{p}(U)=\frac{4}{9}\left(L(U)+\left[L(US)-L(S)\right]\right),$ (43)
where $L(U)$, $L(US)$, and $L(S)$ are linear entropies of the two-qubit gate $U$, its mirror gate, and SWAP gate respectively.
Gate typicality of a two-qubit gate $U$ was defined as antisymmetric combination of $L(U)$ and $L(US)-L(S)$ [11].
$g_{t}(U)=\frac{4}{3}\left(L(U)-\left[L(US)-L(S)\right]\right).$ (44)
It can be noted that six times the right hand side of Eq. 42 with $e^{ih_{k}}$ replaced by its complex conjugate provides the expression of gate typicality of two-qubit gates. That is,
$g_{t}=\frac{1}{12}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{-ih_{k}}|^{2}.$ (45)
This replacement of $e^{ih_{k}}$ by its complex conjugate in Eq. 42 can be geometrically interpreted as follows: All the three points $e^{ih_{j}}$, $e^{ih_{k}}$, and $e^{-ih_{k}}$ in the argand diagram define a triangle as shown in Fig. 4. The side connecting $e^{ih_{k}}$ and $e^{-ih_{k}}$ is parallel to the imaginary axis. The side connecting $e^{ih_{j}}$ and $e^{ih_{k}}$, and the side connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ are reflection of each other about the median from $e^{ih_{j}}$. We refer to the chord connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ as the median reflected chord corresponding to the chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$. For each chord defining entangling power (shown in Fig. 2a) their corresponding median reflected chords define gate typicality. Thus, replacing the symmetric combination of $L(U)$ and $L(US)-L(S)$ in the definition of entangling power by its antisymmetric combination to define gate typicality corresponds to replacing the squared length of the chords describing the entangling power by the squared length of their corresponding median reflected chords. It has to be noted that if any chord describing the entangling power is parallel to the imaginary axis, then the length of the corresponding chord describing the gate typicality is zero.
Gate typicality can also be described using only three median reflected chords as shown below.
$g_{t}=\frac{1}{6}\sum_{k\neq j}|e^{ih_{j}}-e^{-ih_{k}}|^{2},~{}~{}\textnormal{for}~{}\textnormal{any}~{}j.$ (46)
However, the expression of gate typicality given in Eq. 45 has the same form as the expression of entangling power. Hence, it is useful to describe the difference between the mathematical definitions of entangling power and gate typicality.
Since both the entangling power and gate typicality of a two-qubit gate can be calculated using the chords present in their argand diagram, the linear entropy (Eq. 41) describing the operator entanglement of the two-qubit gate can also be calculated using the chords as follows.
$L=\frac{1}{64}\sum_{j=1}^{3}\sum_{k>j}\left[|e^{ih_{j}}-e^{ih_{k}}|^{2}+2|e^{ih_{j}}-e^{-ih_{k}}|^{2}\right]$ (47)
VI Conclusion
To conclude, the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates are very useful to calculate the nonlocal measures such as entangling power, gate typicality, and operator entanglement of two-qubit gates. In the argand diagram of a two-qubit gate, for each chord describing the entangling power, there exist a median reflected chord describing the gate typicality. We have identified the planes containing the perfect entanglers with chords passing through origin. Among SPEs, only the points representing CNOT and iSWAP equivalence classes are present in more than two planes and all other SPEs are contained in two planes. $\sqrt{\textnormal{SWAP}}$ and $\sqrt{\textnormal{SWAP}}^{\dagger}$ equivalence classes are the only perfect entanglers contained in three planes. Perfect entanglers with a chord passing through origin can transform a pair of orthonormal product states into maximally entangled states. However, not all the perfect entanglers without any chord passing through origin can transform a pair of orthonormal product states into maximally entangled states; only those satisfying one of the conditions derived in this paper can do such a transformation.